OpenStudy (anonymous):

!!!!! MEDALS !!!!! What is the product in scientific notation? (3.5 x 10^-5) (3x10^-10) A. 1.05 x 10^-14 B. 1.05 x 10^-15 C. 1.05 x 10^-13 D. 10.5 x 10^15

OpenStudy (freckles):

first do you know multiplication is commutative so we can rewrite this as: \[(3.5 \cdot 3) \times (10^{-5} \cdot 10^{-10})\]

OpenStudy (freckles):

and after that we might have to adjust just a little to finish writing in scientific notation

OpenStudy (freckles):

can you find the products in the ( )

OpenStudy (anonymous):

10.5 x ( 10^-5 x 10^-10 ) ? I don't understand ( 10^-5 x 10^-10 ) like how? @freckles

OpenStudy (anonymous):

Use exponent rule :D

OpenStudy (freckles):

same base when multiplying means you add the exponents like @science00000 says \[10.5 \times 10^{-5+(-10)} \]

OpenStudy (freckles):

now we have one more step after this to finish up this problem

OpenStudy (anonymous):

\[10^{-5}\times10^{-10}\] When you're multiplying two exponents with same bases, you can simply add the two exponents. In this case, -5 +(-10) = -15, so the answer to \[10^{-5}\times10^{-10}\] would be \[10^{-15}\]

OpenStudy (freckles):

\[10.5 \text{ can be written as } 1.05 \times 10^1 \\ \text{ so you have } 1.05 \times (10^1 \cdot 10^{-5+(-10)})\] do you think you finish this up? :)

OpenStudy (anonymous):

= 1.05 x 10^14 ?

OpenStudy (anonymous):

yes, but 10^-14

OpenStudy (anonymous):

Thank you! :) @freckles @science00000

OpenStudy (anonymous):

No problem :)